OpenStudy (anonymous):

Please help!!! Evan opens a savings account with $5,000. He deposits $75 every month into the account that compounds annually and has a 0.85% interest rate. What will his account total be in 5 years? $8,267.58 $5,597.57 $9,793.30 $12,851.12

OpenStudy (amistre64):

this looks familiar

OpenStudy (amistre64):

5000 + 12(75) is one year, compound it by (1+.0085/12) add 12(75) to that amount for the second year, and compound again do this for 5 years total

OpenStudy (anonymous):

Never got a answer.

OpenStudy (amistre64):

y1: (5000 + 12(75))(1+.0085/12) y2: ((5000 + 12(75))(1+.0085/12) + 12(75))(1+.0085/12) y3: (((5000 + 12(75))(1+.0085/12) + 12(75))(1+.0085/12)+12(75))(1+.0085/12) + .... y1: (Bo + D)k y2: Bok^2 + Dk + Dk^2 y3: Bok^3 + Dk + Dk^2 + Dk^3 yn: Bok^n + Dk + Dk^2 + Dk^3 + Dk^(n-1) Bok^n + Dk (1-k^(n-1))/(1-k) sounds about right to me, but id have to recheck that sum for Dk to be sure

OpenStudy (amistre64):

S = k + k^2 + k^3 + ... + k^n -kS = - k^2 - k^3 - ... -k^n - k^(n+1) ----------------------------------- (1-k)S = k - k^(n+1) S = (k - k^(n+1))/(1-k) S = k (1 - k^n)/(1-k)

OpenStudy (amistre64):

every way i do it i get about 9527 ....

OpenStudy (anonymous):

So ill go with the third choice.

OpenStudy (amistre64):

it does seem the most reasonable to me as well. i might just be reading the information a little wrong is all. OR they worked it with an approximation